Suppose the short run production function can be represented by

Q=60,000L^2-1000L^3 then, determine

A) the level labor employement that maximazes the level of output

B) the level of employment that maximazes APL and the maximum APL


1
Expert's answer
2021-12-30T15:20:24-0500

A) the level labor employment that maximizes the level of output


The total output will be maximized when the marginal product is equal to zero.

"MPL=\dfrac{dQ}{dL}\\[0.3cm]
MPL = 120,000L– 3000L^2=0\\[0.3cm]
L(120,000– 3000L)=0\\[0.3cm]
120,000– 3000L=0\\[0.3cm]
3000L=120,0000\\[0.3cm]
L^*=\dfrac{120,000}{3000}=40"


B) the level of employment that maximizes APL and the maximum APL


The average product is maximized when the marginal product is equal to the average product.


The average product is;

"AP=\dfrac{Q}{L}\\[0.3cm]
AP=60,000L-1000L^2"

Therefore;

"60,000L-1000L^2=120,000L-3000L^2\\[0.3cm]
4000L^2=60000L\\[0.3cm]
4000L=60000\\[0.3cm]
L=\dfrac{60000}{4000}=15"

The maximum average product is equal to;

"AP=60,000(15)-1000(15)^2\\[0.3cm]
AP=675,000"